The FBI wants to determine the effectiveness of their 10 Most Wanted list. To do so, they need to find out the fraction of people who appear on the list that are actually caught.

Step 1 of 2 :

Suppose a sample of 434 suspected criminals is drawn. Of these people, 173 were captured. Using the data, estimate the proportion of people who were caught after being on the 10 Most Wanted list. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

Solution :

Solution :

Given that,

n = 434

x = 173

The proportion of people who were caught after being on the 10 Most Wanted list = x / n = 173 / 434 = 0.399